Part 3

If the country could do both (i) and (ii), would it choose to do so? Why or why not? Again, give numerical answers here to prove your answer. Think about the intuition here as well!

i. A $4 / ream tax on domestic paper production while remaining closed to free trade in paper
ii. No tax on domestic paper production, but allowing free trade in paper with a world price of $4 / ream of paper

If the tax IS enacted AND free trade IS allowed, total social surplus would be:

A. $12, B. $24, C. $16, D. $32

This total social surplus would be represented by area(s):

A. A + B + C + E + F + G + H + I

B. A + B + C + E + H + G + F + K + L

C. A + B + H + C + G + I

D. A + B + H + E + F + K + L

The externality costs resulting from this policy would be represented by area(s):

A.  J + E + F

B. I

C. None of these

D. C + G + I- MSc SDOM. ------- -----Pa - 4 6 8 12 alvms.) [Note: Areas B & G are trapezoids! All other lettered areas are triangles!!
Answer #1

Total social surplus = Area of triangle A + Area of square (B + C + H G) + Area of triangle I

= 1/2 x base x height + Length x Breadth + 1/2 x base x height

= 1/2 x 4 x (12 - 8) + 4 x (8 - 4) + 1/2 x 4 x (4 - 0) = 2 x 4 + 16 + 2 x 4 = 8 + 16 + 8 = $ 32

Answer is D. $ 32

Total Social Surplus = A + B + H + C + G + I

Answer is C. A + B + H + C + G + I

Externality cost is the area of triangle J. So, answer is C) None of these